# Multimodal Laryngoscopic Video Analysis for Assisted Diagnosis of Vocal Fold Paralysis

### v3

@@ 7.1 Detection of Vocal Fold Paralysis
    Impact of metric-enhancing modules: Evaluating the contributions of the Quadratic Fitting (QF) and Diffusion Refinement (DR) modules in improving metric extraction, and thus improving the VFP detection performance.
@@
Accurate midline prediction is critical for analyzing left and right vocal fold movements, as it serves as a reference for measuring angle deviations and calculating distances between sample points on the vocal folds and the midline. However, in real-world applications, the laryngeal camera’s position is not always fixed, complicating metric computation. Calibrating the midline by fitting a quadratic function has proven effective for improving midline prediction, leading to better classification performance. The associated results are detailed in Table 7. The system’s native setting serves as a foundational benchmark from which incremental enhancements are realized. The introduction of the proposed QF method improves the specificity from 63.73% to 67.37%, reducing the risk of misdiagnosing healthy patients. Further improvements are achieved by incorporating DR into the glottis segmentation process. As indicated in Table 7, all the metrics show great improvement, except precision and specificity show a slight decrease. ROC-AUC, accuracy, and F-score improve by approximately 2% in absolute terms. Particularly, sensitivity grows from 85.03% to 90.06%, significantly reducing the probability of missing a negative sample. DR is applied only to images where the glottis is not detected by the YOLO-v5 detector (discussed in Section 2.3), as this technique specifically targets false alarms (FAs) in glottis detection (described in Section 6.4). By reducing false alarms, the overall performance of downstream VFP detection improves.
@@
Figure 9: Confusion matrix of the proposed system for the overall subject-level UVFP classification on the SYSU-A test dataset.
@@ 7.2 Detection of Unilateral Vocal Fold Paralysis
Through the analysis of variance in LVFDyn and RVFDyn, our approach can differentiate between left and right VFP within laryngeal stroboscopic videos. The performance of our detection system for UVFP is initially demonstrated on the SYSU-A dataset, where it effectively distinguishes between left and right VFP instances. Subsequently, the system’s broader performance is evaluated across the entire dataset with the VFP detection results from the first-stage multimodal VFP analysis, providing a comprehensive view of MLVAS’s real-world clinical applicability.
@@ 7.3 Auxiliary Visual Analysis for UVFP
In this subsection, the potential of the novel feature is demonstrated through expert visual analysis. Figure 10 showcases the GAW and VFDyn plots for a patient diagnosed with right VFP from the SYSU-A dataset. The MLVAS system extracts several video highlights, yielding multiple analytical outcomes. The left three figures present the time-normalized GAW, illustrating the global vibratory behavior of the vocal folds during phonation. However, these plots are not capable of differentiating between the movements of the individual vocal folds, thus clinicians are unable to ascertain the paralyzed vocal fold based solely on the GAW. In contrast, the VFDyn analysis, depicted in the figures listed vertically on the right, offers a disambiguated view of each vocal fold’s motion. This granular perspective significantly augments the utility of the metrics provided by the MLVAS for clinical assessment of UVFP. In this specific example, the VFDyn plots display a markedly greater level of oscillation for the LVFDyn in comparison to the RVFDyn, which is consistent with the patient’s right VFP diagnosis, thus validating the clinical utility of the MLVAS-derived metrics.
@@ 8 Conclusion
In this work, we present the Multimodal Laryngoscopic Video Analyzing System (MLVAS), a novel framework that integrates audio and image processing for the advanced diagnosis of laryngeal disorders. The front-end utilizes an audio-based KWS model to capture the full phonation cycle, along with a vocal fold detection module to confirm the presence of vocal folds. Leveraging HSV color space analysis, MLVAS extracts stroboscopic sequences, enabling the automatic generation of video highlights from raw laryngoscopic footage, thereby eliminating the need for manual pre-editing. The back-end integrates both audio and visual modalities for VFP detection, marking the first application of pre-trained audio models for laryngeal disease diagnosis to encode robust audio features. Visual metrics are computed using a two-stage segmentation pipeline, combining U-Net-based segmentation with diffusion refinement to reduce false positives, followed by quadratic midline fitting to derive LVFDyn and RVFDyn, representing the dynamic angle deviations of the left and right vocal folds. The novel visual metrics enable MLVAS not only to detect VFP but also to distinguish left- from right-sided VFP. By uniting audio and visual features in a cohesive framework, MLVAS achieves superior performance on the clinical dataset and provides interpretable visualizations, such as the GAW and VFDyn, equipping clinicians with reliable tools for precise and efficient UVFP diagnosis.
@@ 9 Acknowledgment
This research is funded by the DKU foundation project ”Interdisciplinary Signal Processing Technologies”. Many thanks for the computational resource provided by the Advanced Computing East China Sub-Center.

### v4

@@ 7.1 Detection of Vocal Fold Paralysis
    Impact of metric-enhancing modules: Evaluating the contributions of the QF and Diffusion Refinement (DR) modules in improving metric extraction, and thus improving the VFP detection performance.
@@
Accurate midline prediction is critical for analyzing left and right vocal fold movements, as it serves as a reference for measuring angle deviations and calculating distances between sample points on the vocal folds and the midline. However, in real-world applications, the laryngeal camera’s position is not always fixed, complicating metric computation. Calibrating the midline by fitting a quadratic function has proven effective for improving midline prediction, leading to better classification performance. The associated results are detailed in Table 8. The system's native setting serves as a foundational benchmark from which incremental enhancements are realized. The introduction of the proposed QF method improves the specificity from 63.73% to 67.37%, reducing the risk of misdiagnosing healthy patients. Further improvements are achieved by incorporating DR into the glottis segmentation process. As indicated in Table 8, all the metrics show great improvement, except precision and specificity show a slight decrease. ROC-AUC, accuracy, and F-score improve by approximately 2% in absolute terms. Particularly, sensitivity grows from 85.03% to 90.06%, significantly reducing the probability of missing a negative sample. DR is applied only to images where the glottis is not detected by the YOLO-v5 detector (discussed in Section 2.3), as this technique specifically targets false alarms (FAs) in glottis detection (described in Section 6.4). By reducing false alarms, the overall performance of downstream VFP detection improves.
@@
Table 9: One-tailed paired $t$-test $p$-values ($\alpha=0.05$) comparing the proposed system against ablated and unimodal baselines across key performance metrics. Asterisks (*) denote statistically significant improvements. DR: Diffusion-based refinement. QF: Quadratic fitting. Audio-only: System using only audio modality without visual features.
Proposed System Precision Specificity F1-score Mean(Sens, Spec)
vs. w/o DR 0.019566 0.011240 0.005401 0.004115
vs. w/o DR or QF 0.027885 0.006049 0.033392 0.009962
vs. Audio-only 0.004002 0.001061 0.003680 0.001114
7.1.4 Statistical Significance
[…]
As reported in Table 9, all $p$-values fall below the one-tailed significance threshold of $\alpha=0.05$, indicating that the proposed system achieves statistically significant improvements over the baselines across all evaluation metrics. The first two comparisons isolate the contributions of our enhancement modules. The significant gains confirm that the proposed refinements—quadratic midline fitting and diffusion-based segmentation—substantially improve feature quality and detection accuracy. The third comparison, labeled ``Audio-only,'' evaluates the added value of incorporating visual features (VFDyn) into the system. The statistically significant improvement further demonstrates the importance of visual information for achieving robust and reliable VFP detection.
@@ 7.1.4 Statistical Significance
Figure 9: Confusion matrix of the proposed system for UVFP classification on the SYSU-A test dataset.
@@ 7.2 Detection of Unilateral Vocal Fold Paralysis
Through the analysis of variance in LVFDyn and RVFDyn, our approach can differentiate between left and right VFP within laryngeal stroboscopic videos. The performance of our detection system for UVFP is initially demonstrated on the SYSU-A dataset, where it effectively distinguishes between left and right VFP instances. Subsequently, the system's broader performance is evaluated across the entire dataset with the VFP detection results from the first-stage multimodal VFP analysis, providing a comprehensive view of MLVAS's real-world clinical applicability.
@@ 7.3 Auxiliary Visual Analysis for UVFP
In this subsection, the potential of the novel feature is demonstrated through expert visual analysis. Figure 10 showcases the GAW and VFDyn plots for a patient diagnosed with right VFP from the SYSU-A dataset. The MLVAS system extracts several video highlights, yielding multiple analytical outcomes. The left three figures present the time-normalized GAW, illustrating the global vibratory behavior of the vocal folds during phonation. However, these plots are not capable of differentiating between the movements of the individual vocal folds, thus clinicians are unable to ascertain the paralyzed vocal fold based solely on the GAW. In contrast, the VFDyn analysis, depicted in the figures listed vertically on the right, offers a disambiguated view of each vocal fold's motion. This granular perspective significantly augments the utility of the metrics provided by the MLVAS for clinical assessment of UVFP. In this specific example, the VFDyn plots display a markedly greater level of oscillation for the LVFDyn in comparison to the RVFDyn, which is consistent with the patient's right VFP diagnosis, thus validating the clinical utility of the MLVAS-derived metrics.
@@ 8 Conclusion
In this work, we present the Multimodal Laryngoscopic Video Analyzing System (MLVAS), a novel framework that integrates audio and image processing for the advanced diagnosis of laryngeal disorders. The front-end utilizes an audio-based KWS model to capture the full phonation cycle, along with a vocal fold detection module to confirm the presence of vocal folds. Leveraging HSV color space analysis, MLVAS extracts stroboscopic sequences, enabling the automatic generation of video highlights from raw laryngoscopic footage, thereby eliminating the need for manual pre-editing. The back-end integrates both audio and visual modalities for VFP detection, marking the first application of pre-trained audio models for laryngeal disease diagnosis to encode robust audio features. Visual metrics are computed using a two-stage segmentation pipeline, combining U-Net-based segmentation with diffusion refinement to reduce false positives, followed by quadratic fitting to derive LVFDyn and RVFDyn, representing the dynamic angle deviations of the left and right vocal folds. The novel visual metrics enable MLVAS not only to detect VFP but also to distinguish left- from right-sided VFP. By uniting audio and visual features in a cohesive framework, MLVAS achieves superior performance on the clinical dataset and provides interpretable visualizations, such as the GAW and VFDyn, equipping clinicians with reliable tools for precise and efficient UVFP diagnosis.
@@ 9 Acknowledgment
This research is funded by the DKU foundation project ``Interdisciplinary Signal Processing Technologies''. Many thanks for the computational resource provided by the Advanced Computing East China Sub-Center.